OpenStudy (anonymous):

Two similar pipes are shown. Similar pipes; left pipe has height 21 feet and radius 18 feet; right pipe has height 28 feet. The small pipe has a radius of 18 feet and a height of 21 feet while the large pipe has a height of 28 feet. What is the volume of the large pipe in terms of π?

OpenStudy (anonymous):

Not mine, but The volume of the smaller pipe is: pi * 18*18 * 21 = 6804 pi. The larger pipe is similar to the smaller one and the scale factor from small to large is 21 to 28 which simplifies to 3/4. If two solids are similar the cube of the scale factor is the ratio of the volumes. (3/4) ^3 = 6804 pi / V where V is the volume of the larger solid 27/64 = 6804 pi / V V = (64 * 6804 pi) / 27 = 16 128 pi --> ANS

OpenStudy (anonymous):

similar r8 18/21=r/28 solve it u get r nd den calculate vol